# rowsep=0pt

### v1

@@ Abstract
Automated radiology report generation (R2Gen) has advanced significantly, introducing challenges in accurate evaluation due to its complexity. Traditional metrics often fall short by relying on rigid word-matching or focusing only on pathological entities, leading to inconsistencies with human assessments. To bridge this gap, we introduce ER2Score, an automatic evaluation metric designed specifically for R2Gen. Our metric utilizes a reward model, guided by our margin-based reward enforcement loss, along with a tailored training data design that enables customization of evaluation criteria to suit user-defined needs. It not only scores reports according to user-specified criteria but also provides detailed sub-scores, enhancing interpretability and allowing users to adjust the criteria between different aspects of reports. Leveraging GPT-4, we designed an easy-to-use data generation pipeline, enabling us to produce extensive training data based on two distinct scoring systems, each containing reports of varying quality along with corresponding scores. These GPT-generated reports are then paired as accepted and rejected samples through our pairing rule to train an LLM towards our fine-grained reward model, which assigns higher rewards to the report with high quality. Our reward-control loss enables this model to simultaneously output multiple individual rewards corresponding to the number of evaluation criteria, with their summation as our final ER2Score. Our experiments demonstrate ER2Score’s heightened correlation with human judgments and superior performance in model selection compared to traditional metrics. Notably, our model provides both an overall score and individual scores for each evaluation item, enhancing interpretability. We also demonstrate its flexible training across various evaluation systems.
@@ 1 Introduction
To drive progress in this field, this study proposes ER2Score, an innovative metric tailored specifically for evaluating automated radiology report generation. Leveraging GPT-4’s human-like scoring capacity [11, 12], our method autonomously produces evaluation samples that mimic human judgment. These samples are subsequently utilized to train an LLM-based reward model for automated scoring. In comparison to traditional evaluation metrics, ER2Score substantially improves the alignment with human assessments, leading to a more precise evaluation of report quality. Moreover, instead of merely providing an overall score, our model simultaneously outputs the scores for individual evaluation criteria, improving the interpretability of the assessment results. For example, by combining sub-criteria, we can pinpoint the reasons for a report’s poor quality, e.g., whether due to incorrect lesion location, incorrect severity of findings, or omission of findings. Meanwhile, through LLM-generated samples, our method reduces dependence on costly human annotations, supporting scalable training and adaptability to various criteria. Utilizing the defined criteria, we prompt GPT-4 to generate report samples with varied quality levels, pairing reports of different quality corresponding to the same ground-truth report as "accepted" and "rejected" samples with score margins. These paired samples were then used to fine-tune the pretrained Llama3 model [13] using reward modeling techniques. Our proposed loss function enables this model to produce multiple individual rewards concurrently, each corresponding to one evaluation criteria, which are then summed to produce our final ER2Score. Validating our model on two datasets paired with human evaluations, we found ER2Score aligns more closely with human judgment than other traditional metrics and exhibits versatility to accommodate different evaluation criteria.
@@
Our ER2Score metric offers key advantages over existing approaches. Non-trainable metrics, including NLG-based and clinically relevant ones, correlate poorly with human assessments and cannot adapt to customized criteria. Among trainable metrics, RadCliQ combines non-trainable scores linearly, offering limited improvement and flexibility. Some LLM-based metrics (e.g., G_Rad [14] and FineRadScore [15]), relying on online LLMs, present privacy concerns. The most similar methods, MRScore [12] and GREEN [16], fall short in different ways: MRScore only provides an overall score, while GREEN adds interpretability through free-text analysis but sacrifices scoring accuracy due to the task complexity. Furthermore, GREEN’s fine-tuning of LLMs lacks a dedicated loss function as ours, limiting its sensitivity to nuanced quality differences. Our method achieves a higher Kendall’s Tau correlation with human ratings (0.75 vs. 0.64 for GREEN), while reducing training costs (1x NVIDIA A6000 vs. 8x NVIDIA A100) and inference time. To date, no other trainable metrics demonstrate adaptability to different evaluation criteria as ours.
@@
    Our study presents a novel approach to training LLMs to generate ER2Score, a human-consistent metric designed for automated radiology report evaluation. Through our novel loss function discerning report rankings, we finetune LLMs to produce rewards aligned with our scoring system in a fine-grained manner, enhancing alignment with human evaluations and bolstering assessment accuracy.
@@
    By facilitating a tailored analysis of report components, our ER2Score allows users to customize the evaluation framework to suit their specific needs. This level of customization could contribute to more targeted improvements in report generation. This capacity of ER2Score has been demonstrated by its versatility to accommodate two distinct sets of evaluation standards, respectively.
@@ 3 Method
Figure 1: Overview of ER2Score. The upper portion illustrates the training data generation process, while the lower portion represents the training process for the reward model using LoRA. In the lower portion, the solid line indicates the training phase while the dashed line indicates the inference phase.
@@ 3.2 LLM-based Reward Model
ER2Score is our innovative evaluation metric designed to be versatile across various evaluation frameworks. This LLM-based reward model leverages a pretrained language model, such as Llama3 [21], fine-tuning it to align with human evaluations using pairs of reports guided by our novel reward system. The core of ER2Score is its training process, which involves pairs of reports generated from the same ground-truth report but with different qualities. This pairing mechanism is essential for calibrating the model to distinguish between different quality levels effectively. During training, the model learns to assign higher rewards to high-quality reports while simultaneously generating multiple individual criterion scores. These criterion scores are critical as they provide detailed insights into specific aspects of the report’s quality. At the inference stage, the model predicts rewards for each individual criterion. These rewards are then summed to generate the final ER2Score. To ensure precise differentiation, we also introduce a scoring margin for each criterion and the overall score. This margin enables the model to recognize and learn subtle differences in report quality, enhancing its evaluative capability.
@@
LLM Model.  Our reward model, based on the Llama3 [13] backbone, incorporates a multi-reward head to generate the ER2Score. Llama3 was selected for its exceptional language comprehension with just 6.8M trainable parameters over 7 billion in total. The multi-reward head is a linear projection layer mapping Llama-3’s last layer feature map to an $N\times 1$ vector, where $N$ is the total number of sub-scores. This model is fine-tuned using Low-Rank Adaptation (LoRA) [22] for parameter-efficient fine-tuning (PEFT), allowing effective fine-tuning with minimal parameter changes. Training pairs of "accepted" and "rejected" reports calibrate the model for reward prediction. During training, the model learns to distinguish report qualities by adhering to a scoring margin reflecting quality differences. Sub-scores discern quality differences per report aspect, with their summation producing the overall assessment for generated reports.
@@ 4.1 Datasets
We evaluated the effectiveness of ER2Score by assessing its alignment with expert radiologist evaluations, ensuring that its predictions correlate closely with those of human experts. Our evaluation involved two datasets, ReXVal [23] and Rad-100, each based on a distinct scoring system as described in Section 3.1. This approach allowed us to validate ER2Score across different evaluative standards, exhibiting the model’s adaptability to diverse assessment systems.
@@
Rad-100 Dataset which we developed using the MRScore scoring system, consists of 100 diagnostic reports generated by the conventional R2Gen models. Each report displays varying qualities when compared to its corresponding ground-truth report, which has been randomly sampled from the MIMIC-CXR dataset. Employing this scoring system, an experienced radiologist performs detailed evaluations of each report, assessing both overall performance and individual criteria. These evaluations provide a robust foundation for validating our ER2Score. [^2] The Rad-100 dataset is entirely distinct from the datasets used for training our reward model.
@@ 4.2 Performance on ReXVal Dataset
Correlation Analysis of Sub-criteria.  Table 1 provides a quantitative evaluation of ER2Score on the ReXVal dataset, specifically constructed based on the RadCliQ Scoring System. This assessment highlights significant alignment between ER2Score evaluations and expert radiologist judgments across various error categories, using Kendall’s Tau and Spearman Correlation coefficients as metrics. Notably, the high correlation scores in categories such as “False prediction of a finding” (Kendall’s Tau: 0.680, Spearman: 0.842) and “Omission of a finding” (Kendall’s Tau: 0.507, Spearman: 0.673) demonstrate ER2Score’s capability in accurately identifying common radiological errors, indicating its effectiveness in recognizing significant or typical lesions. Although ER2Score demonstrates strong correlations across most sub-criteria, there are areas for improvement. For example, the scores for “Incorrect location or position of a finding” (Kendall’s Tau: 0.246, Spearman: 0.327) are relatively low, possibly because location and position details are often subtle and challenging to capture accurately. It is worth noting that this also highlights the advantage of ER2Score over methods that provide only an overall score [9, 19, 7]. By providing scores for each sub-criterion, ER2Score allows us to clearly identify specific areas where the model can be enhanced. The statistical significance of the results is underscored by extremely low p-values across all categories, reinforcing the robustness of the correlation between ER2Score and expert evaluations. The overall high scores—0.751 for Kendall’s Tau and 0.910 for Spearman Correlation—further validate the reliability of ER2Score as an evaluation tool, highlighting its potential utility in clinical and research settings for assessing radiology reports.
Table 1: Human Correlations of ER2Score on ReXVal Dataset using RadCliQ scoring criteria.
@@
Comparison with other metrics.  Table 2 compares the performance of different metrics using Kendall’s Tau and Spearman correlation on ReXVal Dataset. The comparison is based on the total score. Unlike ER2Score, the existing metrics cannot be customized to user-specific sub-criteria, making sub-score comparison impossible [^3] GREEN only provides error counts for each subcategory without human correlation, making direct comparison of subscore correlations unfeasible. Additionally, its overall correlation with human assessments is significantly lower than ours..
We evaluate our ER2Score against various NLG metrics, including BLEU-4 [3], ROUGE-L [17], METEOR [18], and CIDEr [24], as well as clinical metrics like BERTScore [19] and RadGraph F1 [7]. We also compare with RadCliQ-based metrics [9] derived from human-annotated error scores, as well as the LLM-based GREEN score.
The table demonstrates that ER2Score exhibits a strong alignment with human judgments, as evidenced by its Kendall’s Tau value of 0.751 and Spearman correlation of 0.910, both surpassing all other evaluated metrics. For instance, traditional NLG metrics like BLEU-4, ROUGE-L, and METEOR show lower correlations, with BLEU-4 achieving a Kendall’s Tau of 0.345 and a Spearman correlation of 0.475. Similarly, clinical metrics such as BERTScore and RadGraph F1, while performing better than traditional NLG metrics, still fall short compared to ER2Score. RadCliQ-v1 metric shows high correlation values, with a Kendall’s Tau of 0.631 and a Spearman correlation of 0.816, indicating its effectiveness in aligning with human evaluations. The GREEN score shows higher correlation values, with Kendall’s Tau 0.640, lower than our 0.751. GREEN requires 1.06 seconds to infer one sample, while our model only needs 0.04 seconds. Notably, GREEN incurs much higher training cost by using 8 NVIDIA A100 GPUs (40GB VRAM) with a batch size of 2,048 for 12 epochs, whereas our model was trained with only 1 A6000 GPU (40GB VRAM) with a batch size of 6 for 4 epochs. However, our ER2Score outperforms all these metrics, highlighting its superior ability to capture the nuances of radiology report generation as judged by experts.
@@ 4.6 Qualitative Analysis
A visual example is provided in Figure 3, demonstrating how the ER2Score correlates with human ratings using the RadCliQ scoring system. As shown, the generated report inaccurately describes the severity of the “left pleural effusion” (highlighted in red), resulting in a high ER2Score for “incorrect severity of a finding”, which aligns with the human rating. Additionally, the report erroneously mentions a “right pleural effusion”, leading to an “incorrect location/position of a finding”, again perceived similarly by both the ER2Score and human ratings. Lastly, the generated report fails to mention the “left retrocardiac opacification”, leading to a score of ‘1.0’ for “false prediction of a finding” from both the ER2Score and the human rating.
Figure 3: An visual example of ER2Score from ReXVal Dataset. The highlighted sentences in reports and their corresponding scores share the same colors.
@@ 5 Conclusions
ER2Score offers a human-correlated and explainable metric for evaluating radiology reports. It allows for more fine-grained scoring, aligning each item of the evaluation rule with its respective sub-score, therefore enhancing the interpretability of assessment results. Leveraging GPT-4’s human-like scoring capacity, we have tailored extensive training samples to fine-tune LLMs toward discerning report qualities using our designed reward loss. Our metric’s adaptability allows for accommodating various scoring criteria.
@@ 6 Ethics Statement
Our ER2Score model, which fine-tunes LLAMA-3 as a reward system, operates entirely locally once trained, eliminating the need for any interactions with GPT-4 during inference. This local deployment ensures that there is no risk of information leakage. GPT-4 is only used to generate training data from MIMIC-CXR dataset. MIMIC-CXR is a public dataset, which has been anonymized and de-identified. The platform Azure OpenAI is HIPAA compliant and ensures the privacy and compliance of medical data (e.g., the data are not accessible to OpenAI).

### v2

@@ Abstract
Automated radiology report generation (R2Gen) has advanced significantly, introducing challenges in accurate evaluation due to its complexity. Traditional metrics often fall short by relying on rigid word-matching or focusing only on pathological entities, leading to inconsistencies with human assessments. To bridge this gap, we introduce ReFINE, an automatic evaluation metric designed specifically for R2Gen. Our metric utilizes a reward model, guided by our margin-based reward enforcement loss, along with a tailored training data design that enables customization of evaluation criteria to suit user-defined needs. It not only scores reports according to user-specified criteria but also provides detailed sub-scores, enhancing interpretability and allowing users to adjust the criteria between different aspects of reports. Leveraging GPT-4, we designed an easy-to-use data generation pipeline, enabling us to produce extensive training data based on two distinct scoring systems, each containing reports of varying quality along with corresponding scores. These GPT-generated reports are then paired as accepted and rejected samples through our pairing rule to train an LLM towards our fine-grained reward model, which assigns higher rewards to the report with high quality. Our reward-control loss enables this model to simultaneously output multiple individual rewards corresponding to the number of evaluation criteria, with their summation as our final ReFINE. Our experiments demonstrate ReFINE’s heightened correlation with human judgments and superior performance in model selection compared to traditional metrics. Notably, our model provides both an overall score and individual scores for each evaluation item, enhancing interpretability. We also demonstrate its flexible training across various evaluation systems.
@@ 1 Introduction
To drive progress in this field, this study proposes ReFINE, an innovative metric tailored specifically for evaluating automated radiology report generation. Leveraging GPT-4’s human-like scoring capacity [11, 12], our method autonomously produces evaluation samples that mimic human judgment. These samples are subsequently utilized to train an LLM-based reward model for automated scoring. In comparison to traditional evaluation metrics, ReFINE substantially improves the alignment with human assessments, leading to a more precise evaluation of report quality. Moreover, instead of merely providing an overall score, our model simultaneously outputs the scores for individual evaluation criteria, improving the interpretability of the assessment results. For example, by combining sub-criteria, we can pinpoint the reasons for a report’s poor quality, e.g., whether due to incorrect lesion location, incorrect severity of findings, or omission of findings. Meanwhile, through LLM-generated samples, our method reduces dependence on costly human annotations, supporting scalable training and adaptability to various criteria. Utilizing the defined criteria, we prompt GPT-4 to generate report samples with varied quality levels, pairing reports of different quality corresponding to the same ground-truth report as "accepted" and "rejected" samples with score margins. These paired samples were then used to fine-tune the pretrained Llama3 model [13] using reward modeling techniques. Our proposed loss function enables this model to produce multiple individual rewards concurrently, each corresponding to one evaluation criteria, which are then summed to produce our final ReFINE. Validating our model on two datasets paired with human evaluations, we found ReFINE aligns more closely with human judgment than other traditional metrics and exhibits versatility to accommodate different evaluation criteria.
@@
Our ReFINE metric offers key advantages over existing approaches. Non-trainable metrics, including NLG-based and clinically relevant ones, correlate poorly with human assessments and cannot adapt to customized criteria. Among trainable metrics, RadCliQ combines non-trainable scores linearly, offering limited improvement and flexibility. Some LLM-based metrics (e.g., G_Rad [14] and FineRadScore [15]), relying on online LLMs, present privacy concerns. The most similar methods, MRScore [12] and GREEN [16], fall short in different ways: MRScore only provides an overall score, while GREEN adds interpretability through free-text analysis but sacrifices scoring accuracy due to the task complexity. Furthermore, GREEN’s fine-tuning of LLMs lacks a dedicated loss function as ours, limiting its sensitivity to nuanced quality differences. Our method achieves a higher Kendall’s Tau correlation with human ratings (0.75 vs. 0.64 for GREEN), while reducing training costs (1x NVIDIA A6000 vs. 8x NVIDIA A100) and inference time. To date, no other trainable metrics demonstrate adaptability to different evaluation criteria as ours.
@@
    Our study presents a novel approach to training LLMs to generate ReFINE, a human-consistent metric designed for automated radiology report evaluation. Through our novel loss function discerning report rankings, we finetune LLMs to produce rewards aligned with our scoring system in a fine-grained manner, enhancing alignment with human evaluations and bolstering assessment accuracy.
@@
    By facilitating a tailored analysis of report components, our ReFINE allows users to customize the evaluation framework to suit their specific needs. This level of customization could contribute to more targeted improvements in report generation. This capacity of ReFINE has been demonstrated by its versatility to accommodate two distinct sets of evaluation standards, respectively.
@@ 3 Method
Figure 1: Overview of ReFINE. The upper portion illustrates the training data generation process, while the lower portion represents the training process for the reward model using LoRA. In the lower portion, the solid line indicates the training phase while the dashed line indicates the inference phase.
@@ 3.2 LLM-based Reward Model
ReFINE is our innovative evaluation metric designed to be versatile across various evaluation frameworks. This LLM-based reward model leverages a pretrained language model, such as Llama3 [21], fine-tuning it to align with human evaluations using pairs of reports guided by our novel reward system. The core of ReFINE is its training process, which involves pairs of reports generated from the same ground-truth report but with different qualities. This pairing mechanism is essential for calibrating the model to distinguish between different quality levels effectively. During training, the model learns to assign higher rewards to high-quality reports while simultaneously generating multiple individual criterion scores. These criterion scores are critical as they provide detailed insights into specific aspects of the report’s quality. At the inference stage, the model predicts rewards for each individual criterion. These rewards are then summed to generate the final ReFINE. To ensure precise differentiation, we also introduce a scoring margin for each criterion and the overall score. This margin enables the model to recognize and learn subtle differences in report quality, enhancing its evaluative capability.
@@
LLM Model.  Our reward model, based on the Llama3 [13] backbone, incorporates a multi-reward head to generate the ReFINE. Llama3 was selected for its exceptional language comprehension with just 6.8M trainable parameters over 7 billion in total. The multi-reward head is a linear projection layer mapping Llama-3’s last layer feature map to an $N\times 1$ vector, where $N$ is the total number of sub-scores. This model is fine-tuned using Low-Rank Adaptation (LoRA) [22] for parameter-efficient fine-tuning (PEFT), allowing effective fine-tuning with minimal parameter changes. Training pairs of "accepted" and "rejected" reports calibrate the model for reward prediction. During training, the model learns to distinguish report qualities by adhering to a scoring margin reflecting quality differences. Sub-scores discern quality differences per report aspect, with their summation producing the overall assessment for generated reports.
@@ 4.1 Datasets
We evaluated the effectiveness of ReFINE by assessing its alignment with expert radiologist evaluations, ensuring that its predictions correlate closely with those of human experts. Our evaluation involved two datasets, ReXVal [23] and Rad-100, each based on a distinct scoring system as described in Section 3.1. This approach allowed us to validate ReFINE across different evaluative standards, exhibiting the model’s adaptability to diverse assessment systems.
@@
Rad-100 Dataset which we developed using the MRScore scoring system, consists of 100 diagnostic reports generated by the conventional R2Gen models. Each report displays varying qualities when compared to its corresponding ground-truth report, which has been randomly sampled from the MIMIC-CXR dataset. Employing this scoring system, an experienced radiologist performs detailed evaluations of each report, assessing both overall performance and individual criteria. These evaluations provide a robust foundation for validating our ReFINE. [^2] The Rad-100 dataset is entirely distinct from the datasets used for training our reward model.
@@ 4.2 Performance on ReXVal Dataset
Correlation Analysis of Sub-criteria.  Table 1 provides a quantitative evaluation of ReFINE on the ReXVal dataset, specifically constructed based on the RadCliQ Scoring System. This assessment highlights significant alignment between ReFINE evaluations and expert radiologist judgments across various error categories, using Kendall’s Tau and Spearman Correlation coefficients as metrics. Notably, the high correlation scores in categories such as “False prediction of a finding” (Kendall’s Tau: 0.680, Spearman: 0.842) and “Omission of a finding” (Kendall’s Tau: 0.507, Spearman: 0.673) demonstrate ReFINE’s capability in accurately identifying common radiological errors, indicating its effectiveness in recognizing significant or typical lesions. Although ReFINE demonstrates strong correlations across most sub-criteria, there are areas for improvement. For example, the scores for “Incorrect location or position of a finding” (Kendall’s Tau: 0.246, Spearman: 0.327) are relatively low, possibly because location and position details are often subtle and challenging to capture accurately. It is worth noting that this also highlights the advantage of ReFINE over methods that provide only an overall score [9, 19, 7]. By providing scores for each sub-criterion, ReFINE allows us to clearly identify specific areas where the model can be enhanced. The statistical significance of the results is underscored by extremely low p-values across all categories, reinforcing the robustness of the correlation between ReFINE and expert evaluations. The overall high scores—0.751 for Kendall’s Tau and 0.910 for Spearman Correlation—further validate the reliability of ReFINE as an evaluation tool, highlighting its potential utility in clinical and research settings for assessing radiology reports.
Table 1: Human Correlations of ReFINE on ReXVal Dataset using RadCliQ scoring criteria.
@@
Comparison with other metrics.  Table 2 compares the performance of different metrics using Kendall’s Tau and Spearman correlation on ReXVal Dataset. The comparison is based on the total score. Unlike ReFINE, the existing metrics cannot be customized to user-specific sub-criteria, making sub-score comparison impossible [^3] GREEN only provides error counts for each subcategory without human correlation, making direct comparison of subscore correlations unfeasible. Additionally, its overall correlation with human assessments is significantly lower than ours..
We evaluate our ReFINE against various NLG metrics, including BLEU-4 [3], ROUGE-L [17], METEOR [18], and CIDEr [24], as well as clinical metrics like BERTScore [19] and RadGraph F1 [7]. We also compare with RadCliQ-based metrics [9] derived from human-annotated error scores, as well as the LLM-based GREEN score.
The table demonstrates that ReFINE exhibits a strong alignment with human judgments, as evidenced by its Kendall’s Tau value of 0.751 and Spearman correlation of 0.910, both surpassing all other evaluated metrics. For instance, traditional NLG metrics like BLEU-4, ROUGE-L, and METEOR show lower correlations, with BLEU-4 achieving a Kendall’s Tau of 0.345 and a Spearman correlation of 0.475. Similarly, clinical metrics such as BERTScore and RadGraph F1, while performing better than traditional NLG metrics, still fall short compared to ReFINE. RadCliQ-v1 metric shows high correlation values, with a Kendall’s Tau of 0.631 and a Spearman correlation of 0.816, indicating its effectiveness in aligning with human evaluations. The GREEN score shows higher correlation values, with Kendall’s Tau 0.640, lower than our 0.751. GREEN requires 1.06 seconds to infer one sample, while our model only needs 0.04 seconds. Notably, GREEN incurs much higher training cost by using 8 NVIDIA A100 GPUs (40GB VRAM) with a batch size of 2,048 for 12 epochs, whereas our model was trained with only 1 A6000 GPU (40GB VRAM) with a batch size of 6 for 4 epochs. However, our ReFINE outperforms all these metrics, highlighting its superior ability to capture the nuances of radiology report generation as judged by experts.
@@ 4.6 Qualitative Analysis
A visual example is provided in Figure 3, demonstrating how the ReFINE correlates with human ratings using the RadCliQ scoring system. As shown, the generated report inaccurately describes the severity of the “left pleural effusion” (highlighted in red), resulting in a high ReFINE for “incorrect severity of a finding”, which aligns with the human rating. Additionally, the report erroneously mentions a “right pleural effusion”, leading to an “incorrect location/position of a finding”, again perceived similarly by both the ReFINE and human ratings. Lastly, the generated report fails to mention the “left retrocardiac opacification”, leading to a score of ‘1.0’ for “false prediction of a finding” from both the ReFINE and the human rating.
Figure 3: An visual example of ReFINE from ReXVal Dataset. The highlighted sentences in reports and their corresponding scores share the same colors.
@@ 5 Conclusions
ReFINE offers a human-correlated and explainable metric for evaluating radiology reports. It allows for more fine-grained scoring, aligning each item of the evaluation rule with its respective sub-score, therefore enhancing the interpretability of assessment results. Leveraging GPT-4’s human-like scoring capacity, we have tailored extensive training samples to fine-tune LLMs toward discerning report qualities using our designed reward loss. Our metric’s adaptability allows for accommodating various scoring criteria.
@@ 6 Ethics Statement
Our ReFINE model, which fine-tunes LLAMA-3 as a reward system, operates entirely locally once trained, eliminating the need for any interactions with GPT-4 during inference. This local deployment ensures that there is no risk of information leakage. GPT-4 is only used to generate training data from MIMIC-CXR dataset. MIMIC-CXR is a public dataset, which has been anonymized and de-identified. The platform Azure OpenAI is HIPAA compliant and ensures the privacy and compliance of medical data (e.g., the data are not accessible to OpenAI).